# Mixstyle based Domain Generalization for Sound Event Detection with Heterogeneous Training Data

### v1

@@ Abstract
This work explores domain generalization (DG) for sound event detection (SED), advancing adaptability towards real-world scenarios. Our approach employs a mean-teacher framework with domain generalization to integrate heterogeneous training data, while preserving the SED model performance across the datasets. Specifically, we first apply mixstyle to the frequency dimension to adapt the mel-spectrograms from different domains. Next, we use the adaptive residual normalization method to generalize features across multiple domains by applying instance normalization in the frequency dimension. Lastly, we use the sound event bounding boxes method for post-processing. Our approach integrates features from bidirectional encoder representations from audio transformers and a convolutional recurrent neural network. We evaluate the proposed approach on DCASE 2024 Challenge Task 4 dataset, measuring polyphonic SED score (PSDS) on the DESED dataset and macro-average pAUC on the MAESTRO dataset. The results indicate that the proposed DG-based method improves both PSDS and macro-average pAUC compared to the challenge baseline.
Index Terms—  sound event detection, semi-supervised learning, domain generalization, mixstyle
@@ 1 Introduction
Sound event detection (SED) [1, 2] involves identifying and classifying sound events from acoustic signals along with their timestamps across various environments. In the recent years, deep learning models [3, 4, 5, 6, 7, 8] have been witnessed success in SED research. However, these models require large amount of strongly labelled data, which are costly and time-consuming to obtain. To address these issues, weakly supervised or semi-supervised learning techniques are used more frequently and also considered under the latest edition of DCASE 2024 Challenge Task 4 [9]. The current edition utilizes a new dataset namely, MAESTRO Real [10] that is softly labelled together with widely used DESED [11] dataset containing strongly labelled data for encouraging SED model development under heterogeneous training conditions. However, training a robust SED model with heterogeneous training data is challenging due to the fact that the labels may not be consistent across different datasets apart from the domain mismatch due to various differences in the multiple datasets.
@@
Deep neural networks (DNNs) often struggle to generalize to unseen domains due to the domain mismatch mentioned above, leading to poor results in real-world scenarios. Domain generalization (DG) [12] has become an essential research topic across fields such as computer vision, audio processing, and natural language processing. Techniques like adversarial gradients, feature-based augmentation, and style transfer normalization have been used to address device domain mismatch, such as in the DCASE challenge Task 1 [13]. Thus, employing domain generalization for heterogeneous training data from different domains for sound event detection remains underexplored. In previous editions of DCASE Challenge Task 4 [14], researchers have explored solutions for domain mismatch. For example, domain adaptation methods have been used to efficiently exploit synthetic strong-labeled data by using domain classifiers, considering the gap between synthetic and real audio data in the DESED dataset [11].
@@
The DCASE 2024 Challenge Task 4 focuses on development of a single SED model for detection of sound classes in DESED and MAESTRO datasets using their training set collectively, in contrast to the previous edition in 2023 [5], where separate SED model development for the two dataset was required. The goal of the SED model is still to provide event classes along with their time boundaries, even with multiple overlapping events. At the same time this task emphasizes on leveraging training data with varying annotation granularity (temporal resolution, soft/hard labels). Systems evaluated on the basis of labels with different granularity can help to understand their behavior and robustness for various applications. In addition, the target classes in different datasets also differ, so sound labels present in one dataset might not be annotated in another. Therefore, the developed SED system needs to handle potentially missing target labels during training and perform without knowing the origin of the audio clips at evaluation time.
In order to address the domain mismatch issue in DCASE 2024 Challenge Task 4, we propose a novel mixstyle-based DG approach for SED with heterogeneous training data. We first analyze the relationship between the domain and the statistics of each feature dimension. Specifically, we mix styles of training instances in the frequency dimension, resulting in novel domains being synthesized implicitly. This increases the domain diversity of the source domains and thereby enhances the generalizability of the trained model. In addition, we use the adaptive residual normalization (AdaResNorm) method to generalize features across multiple domains by applying instance normalization in the frequency dimension. Further, we use the sound event bounding boxes method for post-processing. With above methods, the developed SED system is expected to perform more effectively even though the training data originates from different sources. We also submitted our developed system based on the proposed method for DCASE 2024 Challenge Task 4 participation [15].
2 Proposed Domain Generalization for SED
2.1 Analyzing domain mismatch in heterogeneous data
Domain mismatch in heterogeneous data is mostly due to the original sources of data or how they were collected. In DCASE 2024 Task 4 Challenge, the audio files in DESED were sampled from public sources, such as Freesound and YouTube, while audio clips in MAESTRO was recorded in real-life scenarios. Although all the sound classes in DESED and MAESTRO datasets are not same, some events in DESED are mapped to similar classes in MAESTRO. For example, in DESED, “speech” is a super-class for “people talking,” “children’s voices,” and “announcements” in MAESTRO. “Dishes” in DESED corresponds to “cutlery and dishes” in MAESTRO, and “dog” in DESED is a super-class for “dog bark”. This mapping ensures the network to behave similarly for loss computation during training in case of these mapped classes irrespective of their original sources. However, this class mapping may cause a domain mismatch because the audio features of these classes might differ significantly between the two datasets.
To analyze these differences, we use a 2D convolutional neural network (CNN) as convolutional layers applied to extract local invariant features. We denote the input time-frequency representations of 2D CNNs as $X\in R^{F\times T}$, where $F$ and $T$ are the numbers of frequency bins and frames, respectively. With a batch size of $N$, we can represent feature maps as $M\in R^{N\times C\times F\times T}$, where $C$ is the number of channels. Following [13], we utilize instance statistics across a specific dimension, i.e., mean and standard deviation (std), to analyze audio characteristics in 2D CNNs.
Specifically, the frequency-wise statistics can be formulated as:
$s^{(F)}=\textrm{Concat}(\mu(F),\sigma(F))$ (1)
$s^{(C)}=\textrm{Concat}(\mu(C),\sigma(C))$ (2)
We compare $s^{(F)}$ and $s^{(C)}$ using 2D t-SNE visualization [16] in Figure 1, where the feature maps are generated by the CRNN baseline model of DCASE 2024 Task 4. It is observed that the features from different domains are better separated with frequency-wise statistics than that with channel-wise statistics, demonstrating that the frequency feature dimension carries more domain-relevant information in comparison to the channel dimension.
@@ 2.2 Frequency-wise MixStyle
We analyzed the relationship between the domain and the statistics of each feature dimension, which revealed that the frequency feature dimension carries more domain-relevant information than the channel dimension. Therefore, we propose to use Freq-MixStyle, which normalizes the frequency bands of spectrograms and then denormalizes them with mixed frequency components from two different recordings. The mixing coefficient specifies the shape of the Beta distribution. More specifically, given an input batch $x$, Freq-MixStyle first generates a reference batch $\hat{x}$ from $x$. When both datasets are given in one single batch, $x$ is sampled from two different domains $d$ (DESED) and $m$ (MAESTRO), e.g., $x=[x_{d},x_{m}]$. Then, $\hat{x}$ is obtained by swapping the position of $x_{d}$ and $x_{m}$ as shown in Figure 2, followed by a shuffling operation along the batch dimension applied to each batch. After shuffling, Freq-MixStyle computes the feature statistics, $(\mu(F,x),\sigma(F,x))$ and $(\mu(F,\hat{x}),\sigma(F,\hat{x}))$. Here, $\mu(F,x)$ and $\sigma(F,x)$ are the mean and standard deviation of each instance from the $x$ across $F$-axis. Freq-MixStyle then generates a mixture of feature statistics by:
$\alpha_{mix}=\lambda\mu(F,x)+(1-\lambda)\mu(F,\hat{x}),$ (3)
$\beta_{mix}=\lambda\sigma(F,x)+(1-\lambda)\sigma(F,\hat{x}),$ (4)
@@
[…]
Freq-MixStyle for domain generalization is beneficial as it helps the SED model learn more robust frequency-wise features by simulating new domains during the training process. This increases the diversity of the training data, allowing the SED model to perform effectively on unseen domains. The effect of Freq-MixStyle is regularized by a second parameter $p$ that controls the probability of Freq-MixStyle applied to a batch of recordings. We will experiment with different values of this probability to explore its effects.
@@ 2.3 Adaptive residual normalization
Figure 2: Illustration of our proposed method. The orange and blue labels denote DESED and MAESTRO dataset, respectively.
While Freq-MixStyle mixes the feature statistics of two instances to create new styles and enhance domain generalization, we also employ adaptive residual normalization (AdaResNorm) to focus on adjusting the normalization process. The AdaResNorm is based on the frequency-instance normalization (FreqIN) which generalizes the features on multiple domains by applying instance normalization in the frequency dimension.
[…]
@@
Here $a$, $b$, and $c$ are trainable parameters for balancing, scaling, and shifting, respectively. By adding trainable parameters to control the trade-off between identity and FreqIN, the normalization behavior can be adaptively adjusted based on the characteristics of the data and the requirements of the domain. This can help mitigate the information loss that might occur during the MixStyle process. Additionally, by adjusting the normalization dynamically based on the data characteristics, AdaResNorm can enhance the robustness of the model when dealing with the diverse and mixed domains created by Freq-MixStyle. Thus, integrating AdaResNorm with Freq-MixStyle provides a more effective domain generalization by leveraging the strengths of both methods. It is noted that the adaptive residual normalization is inserted after the first convolution layer.
@@ 2.4 Sound event bounding box-based post-processing
Inspired by bounding box predictions in image object detection [18], the SEBBs [19] are one-dimensional bounding boxes defined by event onset time, event offset time, sound class, and confidence. They represent sound event candidates with a confidence score. The final SED is derived by applying class-wise event-level thresholding to SEBBs’ confidences. For high sensitivity or recall (few missed hits), a low detection threshold is used to detect events even when the confidence of the system is low. For high precision (few false alarms), a higher threshold detects only events with high confidence. SEBBs allow controlling system sensitivity without affecting the detection of an event’s onset and offset times, unlike frame-level thresholding.
@@
2.5 Training model with DG
The backbone model, inherited from previous DCASE Task 4 challenge editions [5, 14], consists of a CRNN [3] that uses self-supervised features from the pre-trained BEATs [20] model. We integrate the MAESTRO data and DESED data (strong, weak, and unlabeled) in one batch. First, we shuffle and mix the MAESTRO and DESED strong data using Freq-MixStyle. After the first 2D CNN layer, we employ the AdaResNorm. The CRNN includes a 2D CNN encoder with 7 convolutional layers, batch normalization, gated linear units, and dropout, followed by a bi-directional GRU (biGRU) layer. BEATs features are concatenated with CNN features before the BiGRU layer. Average pooling adjusts the BEATs features to match the CNN sequence length. Attention pooling then derives clip-wise and frame-wise posteriors. During training, the BEATs model remains frozen, and the mean-teacher framework [21, 22] leverages unlabeled and weakly labeled data. The attention pooling uses the softmax over classes, with values for unlabeled classes masked to negative infinity before applying softmax. After we output the frame level prediction, we use the cSEBBs to further improve the performance on the DESED dataset.
Table 1: Performance in PSDS and mpAUC of different single-systems on the DESED development set (Dev-PSDS), DESED public evaluation set (PubEval-PSDS), and MAESTRO validation set (mpAUC) including frequency-dynamic convolution (FDY), domain generalization (DG), and adaptive residual normalization (AdaResNorm). The joint score is based on the sum of PubEval-PSDS (cSEBBs) and mpAUC.
System Dev-PSDS PubEval-PSDS (raw) PubEval-PSDS (cSEBBs) mpAUC Joint score Parameters
Baseline 0.493 0.549 - 0.721 1.270 1.8M
Baseline + FDY 0.508 0.596 0.601 0.728 1.329 3.4M
Baseline + MixStyle 0.516 0.573 0.583 0.724 1.307 1.8M
Baseline + FDY + MixStyle 0.520 0.596 0.603 0.737 1.340 3.4M
Baseline + MixStyle + AdaResNorm 0.520 0.574 0.588 0.726 1.314 1.8M
Baseline + FDY + MixStyle + AdaResNorm 0.526 0.598 0.604 0.739 1.343 3.4M
@@ 3.1 Datasets
3.1.1 DESED dataset
The DESED dataset [11] consists of 10-second audio clips recorded in domestic environments or synthesized to mimic such environments. It features sound events from 10 classes: alarm bell ringing, blender, cat, dishes, dog, electric shaver/toothbrush, frying, running water, speech, and vacuum cleaner. The synthetic part of the dataset is generated with Scaper [23], using foreground events from Freesound datasets and backgrounds from YouTube videos and the Freesound subset of the MUSAN [24] dataset. The real-world recordings come from AudioSet [25] and include a weakly annotated set (1,578 clips), an unlabeled set (14,412 clips), and a strongly annotated portion (3,470 clips).
3.1.2 MAESTRO Real dataset
The MAESTRO Real dataset [10] consists of a development set (6,426 clips) and an evaluation set of long-form real-world recordings. This dataset contains multiple temporally strong annotated events with soft labels from 17 classes. However, only 11 classes are evaluated in this challenge because the other 6 do not have confidence levels over 0.5. These evaluated classes are: birds_singing, car, people_talking, footsteps, children_voices, wind_blowing, brakes_squeaking, large_vehicle, cutlery_and_dishes, metro_approaching, and metro_leaving. This data was annotated using crowdsourcing, combining temporally-weak labeling with a sliding window approach to determine events’ temporal localization. To obtain soft labels, annotations from multiple annotators are aggregated via MACE [26]. The recordings are taken from the TUT Acoustic Scenes 2016 dataset [27] and are between 3 to 5 minutes long.
3.2 Implementation details
The audio clips are resampled to 16 kHz mono using Librosa and then segmented with a window size of 2048 samples and a hop length of 256 samples. Short-time Fourier transform and Mel-filters are used to create log-mel spectrograms spanning 0 to 8 kHz. The clips shorter than 10 seconds are padded with silence. A batch size of 60 is used, with the distribution: 1/5 MAESTRO, 1/10 synth, 1/10 synth+strong, 1/5 weak, and 2/5 unlabeled. The training process includes 50 warmup epochs, up to 300 epochs, and an epoch decay of 100. Gradient clipping is set at 5.0, the EMA factor for the mean teacher is 0.999, and validation occurs every 10 epochs. The Adam optimizer with a learning rate of 0.001 is used. An exponential warmup is applied for the first 50 epochs, with no early stopping. The probability $p$ of apply Freq-MixStyle we set 0.5 and coefficient to the beta distribution we choose 0.6 in all experiments.

### v3

# DG-SED: Domain Generalization for Sound Event Detection with Heterogeneous Training Data

@@ Abstract
This work explores domain generalization (DG) for sound event detection (SED), advancing adaptability to real-world scenarios. Our approach employs a mean-teacher framework with domain generalization named DG-SED to integrate heterogeneous training data while preserving the SED model performance across the datasets. Specifically, we first apply mixstyle to the frequency dimension to adapt the mel-spectrograms from different domains. Next, we use the adaptive residual normalization method to generalize features across multiple domains by applying instance normalization in the frequency dimension. Lastly, we use the sound event bounding boxes method for post-processing. We evaluate the proposed approach DG-SED on the DCASE 2024 Challenge Task 4, measuring PSDS on the DESED dataset and macro-average pAUC on the MAESTRO dataset. The results indicate that the proposed DG-SED method improves both PSDS and macro-average pAUC compared to the baselines. The code will be released in due course.
@@ 1 Introduction
Sound event detection (SED) [1] involves identifying and classifying sound events from acoustic signals along with their timestamps across various environments. It is a core task in many audio-processing applications such as home surveillance and urban computing. In recent years, deep learning models [2, 3, 4, 5, 6, 7] have witnessed success in SED research. However, these models require a large amount of strongly labeled data, which is costly and time-consuming to obtain. To address these issues, weakly supervised or semi-supervised learning techniques are used and also considered under the latest edition of DCASE 2024 Challenge Task 4 [8]. The current edition utilizes a new dataset namely, MAESTRO Real [9] that is softly labeled together with the widely used DESED [10] dataset containing strongly labeled data for SED model development under heterogeneous training conditions. However, training a robust SED model with heterogeneous training data is challenging because the labels may not be consistent across different datasets, apart from the domain mismatch caused by differences in the multiple datasets [8].
@@
Deep neural networks (DNNs) often struggle to generalize to unseen domains due to the domain mismatch mentioned above, leading to poor results in real-world applications. Domain generalization (DG) [11, 12, 13] has become an essential research topic across different fields. Techniques like feature-based augmentation, and style transfer normalization have been used to address device domain mismatch, such as in the DCASE challenge Task 1 [14, 15]. Thus, employing domain generalization for heterogeneous training data from different domains for sound event detection remains underexplored. In previous editions of DCASE Challenge Task 4, researchers have explored solutions for domain mismatch. For example, domain adaptation methods have been used to efficiently exploit synthetic strong-labeled data by using domain classifiers, considering the gap between synthetic and real data in DESED [10].
@@
The DCASE 2024 Challenge Task 4 focuses on development of a single SED model for detection of sound classes in DESED and MAESTRO datasets using their training set collectively, in contrast to the previous edition in 2023 [4, 16], where separate SED model development for the two dataset was required. The goal of the SED model is still to provide event classes along with their time boundaries, even with multiple overlapping events. At the same time, this task emphasizes leveraging training data with varying annotation granularity (temporal resolution, soft/hard labels). Systems evaluated based on labels with different granularity can help to understand their behavior and robustness for various applications. In addition, the target classes in different datasets also differ, so sound labels present in one dataset might not be annotated in another. Therefore, the developed SED system needs to handle potentially missing target labels during training and perform without knowing the origin of the audio clips at evaluation time.
@@
Several studies have focused on domain generalization in SED. The Mean Teacher method, is a key semi-supervised learning method for DG in SED [17]. However, it still struggles with distribution mismatch between synthetic and real audio [18]. Domain adaptation, widely used in image and acoustic scene classification [19], offers an alternative solution. Domain adaptation typically requires multiple stages, such as feature extraction, alignment, and mapping between source and target domains. How to achieve domain generalization in SED in a one-stage and end-to-end manner is still underexplored.
In this study, we propose a novel approach named DG-SED for SED with heterogeneous training data to address the domain mismatch issue in DCASE 2024 Challenge Task 4. We first analyze the relationship between the domain and the statistics of each feature dimension. For DG-SED, we first mix styles of training instances in the frequency dimension, resulting in novel domains being synthesized implicitly. This increases the domain diversity of the source domains and thereby enhances the generalizability of the trained model. In addition, we use the adaptive residual normalization (AdaResNorm) module to generalize features across multiple domains by applying instance normalization in the frequency dimension. Further, we use the sound event bounding boxes method for post-processing. With the above modules, the DG-SED system is expected to perform more effectively even though the training data originates from different sources. The code will be released in due course.
2 Proposed DG-SED approach
2.1 Analyzing domain mismatch in heterogeneous data
Domain mismatch in heterogeneous data is mostly due to the sources of data or how they were collected. In DCASE 2024 Task 4 Challenge, the audio files in DESED were sampled from public sources, such as Freesound and YouTube, while audio clips in MAESTRO was recorded in real-life scenarios. Although all the sound classes in DESED and MAESTRO datasets are not same, some events in DESED are mapped to similar classes in MAESTRO. For example, in DESED, “speech” is a super-class for “people talking,” “children’s voices,” and “announcements” in MAESTRO. This mapping ensures the network to behave similarly during training in case of these mapped classes irrespective of their original sources. However, this class mapping may cause a domain mismatch because the audio features of these classes might differ significantly between the two datasets.
To analyze these differences, we use a 2D convolutional neural network (CNN) as convolutional layers which are applied to extract local invariant features. We denote the input time-frequency representations of 2D CNNs as $X\in R^{F\times T}$, where $F$ and $T$ are the numbers of frequency bins and frames, respectively. With a batch size of $N$, we can represent feature maps as $M\in R^{N\times C\times F\times T}$, where $C$ is the number of channels. Following [14], we utilize instance statistics across a specific dimension, i.e., mean and standard deviation (std), to analyze audio characteristics in 2D CNNs. Specifically, the frequency-wise statistics can be formulated as:
$s^{(F)}=\textrm{Concat}(\mu(F),\sigma(F))$ (1)
$s^{(C)}=\textrm{Concat}(\mu(C),\sigma(C))$ (2)
We compare $s^{(F)}$ and $s^{(C)}$ using 2D t-SNE visualization in Figure 1, where the feature maps are generated by the CRNN baseline model of DCASE 2024 Task 4. It is observed that the features from different domains are better separated with frequency-wise statistics than that with channel-wise statistics, demonstrating that the frequency feature dimension carries more domain-relevant information in comparison to the channel dimension.
@@ 2.2 Frequency-wise MixStyle
We analyzed the relationship between the domain and the statistics of each feature dimension in Section 2.1, which revealed that the frequency feature dimension carries more domain-relevant information than the channel dimension. Therefore, we propose to use Freq-MixStyle, which normalizes the frequency bands of spectrograms and then denormalizes them with mixed frequency components from two different recordings. The mixing coefficient specifies the shape of the Beta distribution. More specifically, given an input batch $x$, Freq-MixStyle first generates a reference batch $\hat{x}$ from $x$. When both datasets are given in one single batch, $x$ is sampled from two different domains $d$ (DESED) and $m$ (MAESTRO), e.g., $x=[x_{d},x_{m}]$. Then, $\hat{x}$ is obtained by swapping the position of $x_{d}$ and $x_{m}$ as shown in Figure 2, followed by a shuffling operation along the batch dimension applied to each batch. After shuffling, Freq-MixStyle computes the feature statistics, $(\mu(F,x),\sigma(F,x))$ and $)(\mu(F,\hat{x}),\sigma(F,\hat{x}$)). Here, $\mu(F,x)$ and $\sigma(F,x)$ are the mean and standard deviation of each instance from the $x$ across $F$-axis. Freq-MixStyle then generates a mixture of feature statistics by:
[…]
@@
$\text{Freq-MixStyle}(x,\hat{x})=\alpha_{mix}\odot\frac{x-\mu(F,x)}{\sigma(F,x)}+\beta_{mix}$ (5)
Freq-MixStyle for domain generalization is beneficial as it helps the SED model learn more robust frequency-wise features by simulating new domains during the training process.
@@ 2.3 Adaptive residual normalization
Figure 2: Illustration of our proposed DG-SED method. The orange and blue labels denote the DESED and MAESTRO datasets, respectively.
While Freq-MixStyle mixes feature statistics of two instances to enhance domain generalization, we also employ adaptive residual normalization (AdaResNorm) to focus on adjusting the normalization process. The AdaResNorm is based on the frequency-instance normalization (FreqIN) which generalizes the features on multiple domains by applying instance normalization in the frequency dimension.
$\text{FreqIN(x)}=\frac{x-\mu(F,x)}{\sqrt{\sigma^{2}(F,x)+\epsilon}}$ (6)
@@
Here $a$, $b$, and $c$ are trainable parameters for balancing, scaling, and shifting, respectively. By adding trainable parameters to control the trade-off between identity and FreqIN, the normalization behavior can be adaptively adjusted by the characteristics of the data and the requirements of the domain. This can help mitigate the information loss that might occur during the MixStyle process. Additionally, AdaResNorm can enhance the robustness of the model when dealing with the diverse and mixed domains created by Freq-MixStyle. Thus, integrating AdaResNorm with Freq-MixStyle provides a more effective domain generalization by leveraging the strengths of both methods. It is noted that the adaptive residual normalization is inserted after the first convolution layer.
@@ 2.4 Sound event bounding box-based post-processing
Inspired by bounding box predictions in image object detection [21], the SEBBs were proposed in [22], which are one-dimensional bounding boxes defined by event onset time, event offset time, sound class, and confidence. They represent sound event candidates with a confidence score. The final SED is derived by applying class-wise event-level thresholding to SEBBs’ confidences. For high sensitivity or recall (few missed hits), a low detection threshold is used to detect events even when the confidence of the system is low. For high precision (few false alarms), a higher threshold detects only events with high confidence. SEBBs allow controlling system sensitivity without affecting the detection of an event’s onset and offset times, unlike frame-level thresholding.
@@
Table 1: Performance in PSDS and mpAUC of different single-systems on the DESED development set (Dev-PSDS), DESED public evaluation set (PubEval-PSDS), and MAESTRO validation set (mpAUC) including DG-SED method, and adaptive residual normalization (AdaResNorm). The joint score is based on the sum of PubEval-PSDS (cSEBBs) and mpAUC. (x) means the improvements compared without the DG-SED method.
System Dev-PSDS $\uparrow$ PubEval-PSDS (raw) $\uparrow$ PubEval-PSDS (cSEBBs) $\uparrow$ mpAUC $\uparrow$ Joint score $\uparrow$ Parameters
CRNN (DCASE2024 baseline [8]) 0.493 0.549 - 0.721 1.270 (-) 1.8M
FDY-CRNN [3] 0.508 0.596 0.601 0.728 1.329 (-) 3.4M
CRNN + DG-SED (w/o AdaResNorm) 0.516 0.573 0.583 0.724 1.307 ($\uparrow$ 0.037) 1.8M
FDY-CRNN + DG-SED (w/o AdaResNorm) 0.520 0.596 0.603 0.737 1.340 ($\uparrow$ 0.011) 3.4M
[…]
@@ 3 Experiment Setting
3.1 Implementation details
Our implementation is based on a CRNN [2] from previous DCASE Task 4 challenge editions [4, 23]. It is enhanced with self-supervised features from the pre-trained BEATs [24] model. We begin by shuffling and mixing the MAESTRO and DESED strong data using Freq-MixStyle. Freq-MixStyle is used with a probability of 0.5, and a beta distribution coefficient of 0.6 is set for all experiments. After passing through the first 2D CNN layer, AdaResNorm is applied. The CRNN model includes a 2D CNN encoder with 7 convolutional layers, followed by a bi-directional GRU (biGRU) layer. Then, BEATs features are concatenated with the CNN-extracted features before the biGRU layer. Average pooling is applied to the BEATs features to match the sequence length of the CNN encoder. Attention pooling is used to derive both clip-wise and frame-wise posteriors, and the BEATs model remains frozen during training. A mean-teacher [25, 26] framework is utilized to leverage unlabeled and weakly labeled data, with masked softmax applied for unlabeled classes during attention pooling. We further enhance DESED dataset performance using cSEBBs after frame-level predictions. We also employed FDY-CRNN from [3] to demonstrate the cross-backbone ability of our domain generalization method, which uses frequency adaptive kernels to enforce frequency dependency in 2D convolutions. Compared with the baseline CRNN architecture, we replaced the standard 2D CNN with FDY-convolutional blocks.
For audio preprocessing, clips are resampled to 16 kHz mono, segmented with a window size of 2048 samples and a hop length of 256 samples, and converted into log-mel spectrograms. Clips shorter than 10 seconds are padded with silence. The training uses a batch size of 60, with specific data distribution: 1/5 MAESTRO, 1/10 synthetic, 1/10 synthetic + strong, 1/5 weak, and 2/5 unlabeled. We conduct 50 warmup epochs within a total of 300 epochs, with a learning rate of 0.001 and exponential warmup in the first 50 epochs.
@@ 3.2 Datasets
The DESED dataset [10] comprises 10-second audio clips from domestic environments, including both real recordings and synthetic data designed to mimic these settings. The synthetic clips are generated using Scaper [27], while real-world recordings are sourced from AudioSet [28]. The dataset includes a mix of weakly annotated (1,578 clips), unlabeled (14,412 clips), and strongly annotated data (3,470 clips), providing a diverse range of sound events for detection tasks.
The MAESTRO Real dataset [9] features a development set (6,426 clips) and an evaluation set containing long-form real-world recordings. It includes multiple temporally strong annotated events with soft labels, generated through a combination of crowdsourcing and a sliding window approach [29]. The dataset draws recordings from the TUT Acoustic Scenes 2016 dataset [30], covering various real-life scenarios.
Table 2: Performance comparison of our system with others on the DCASE 2024 Task 4 public evaluation set (PSDS) and development set (mpAUC).
System PSDS $\uparrow$ mpAUC $\uparrow$ Joint score $\uparrow$ Params
Chen_NCUT_task4_3 [31] 0.549 0.697 1.246 17M
Kim_GIST-HanwhaVision_task4_1 [32] 0.610 0.686 1.296 4M
Zhang_BUPT_task4_1 [33] 0.572 0.763 1.335 10M
[…]